# appAppendix References (eccv) Package eccv Warning: Package ‘hyperref’ is loaded with option ‘pagebackref’, which is *not* recommended for camera-ready version

### v2

@@ 4.3 Main Result
Main Result on 1024$\times$1024 FacesHQ. To elaborate on the scalability of our method within the Mamba and Stochastic Interpolant framework, we provide comparisons on a high-resolution dataset (1024$\times$1024 FacesHQ) in Table 3. Our primary comparison is against Bidirectional Mamba, a commonly used solution for applying Mamba to 2D image data [57, 96]. With the aim of investigating Mamba’s scalability in large resolutions up to 1,024, we employ the diffusion model on the latent space of $128\times 128$ with a patch size of 2, resulting in 4,096 tokens. The network is trained on 16 A100 GPUs. Notably, our method demonstrates superior results compared to Bidirectional Mamba. Details regarding loss and FID curves can be found in Appendix 9.3. While constrained by GPU resource limitations, preventing longer training duration, we anticipate consistent outperformance of Bidirectional Mamba with extended training duration.
@@
UCF101 dataset. In Table 4, we present our results on the UCF101 dataset, training all methods using 4 A100 GPUs, with further scalability exploration conducted using 16 A100 GPUs. We mainly compare our method consistantly with Bidirection Mamba [96]. For the choice of the 3D Zigzag Mamba, please refer to Appendix 9.5. For Factorized 3D Zigzag Mamba in video processing, we deploy the sst scheme for factorizing spatial and temporal modeling. This scheme prioritizes spatial information complexity over temporal information, hypothesizing that redundancy exists in the temporal domain. Our results consistently demonstrate the superior performance of our method across various scenarios, underscoring the intricacy and effectiveness of our approach.
Figure 9: Visualization of various resolutions on FacesHQ $1024\times 1024$ and MultiModal-CelebA $512\times 512$. Our generated samples present high fidelity across various resolutions.
Table 4: Video Scan Scheme on UCF101 dataset. Our method outperforms the baseline and can achieve even better results when the training scale is increased.
[…]
Bidirection Mamba [96] -4GPU 256.1 320.2
3D Zigzag Mamba -4GPU 238.1 282.3
Factorized 3D Zigzag Mamba -4GPU 216.1 210.2
Bidirection Mamba [96] -16GPU 146.2 201.1
Factorized 3D Zigzag Mamba -16GPU 121.2 140.1
Visualization. We demonstrate the image visualization of our best results on FacesHQ 1024 and MultiModal-CelebA 512 in Figure 9. For the visualization of videos, please refer to Appendix 9.1. It is evident that the visualization is visually pleasing across various resolutions, indicating the efficacy of our methods.
@@ 5 Conclusion
In this paper, we present the Zigzag Mamba Diffusion Model, developed within the Stochastic Interpolant framework. Our initial focus is on addressing the critical issue of spatial continuity. We then devise a Zigzag Mamba block to better utilize the inductive bias in 2D images. Further, we factorize the 3D Mamba into 2D and 1D Zigzag Mamba to facilitate optimization. We empirically design various ablation studies to examine different factors. This approach allows for a more in-depth exploration of the Stochastic Interpolant theory. We hope our endeavor can inspire further exploration in the Mamba network design.
6 Acknowledgement
We would like to thank Timy Phan, Yunlu Chen, Di Wu, Divin Yan for the extensive proofreading.

### v3

@@ 4.3 Main Result
Main Result on 1024$\times$1024 FacesHQ. To elaborate on the scalability of our method within the Mamba and Stochastic Interpolant framework, we provide comparisons on a high-resolution dataset (1024$\times$1024 FacesHQ) in Table 6. Our primary comparison is against Bidirectional Mamba, a commonly used solution for applying Mamba to 2D image data [70, 123]. With the aim of investigating Mamba’s scalability in large resolutions up to 1,024, we employ the diffusion model on the latent space of $128\times 128$ with a patch size of 2, resulting in 4,096 tokens. The network is trained on 16 A100 GPUs. Notably, our method demonstrates superior results compared to Bidirectional Mamba. Details regarding loss, FID curves, and visualization can be found in the Appendix. While constrained by GPU resource limitations, preventing longer training duration, we anticipate consistent outperformance of Bidirectional Mamba with extended training duration.
@@
UCF101 dataset. In Table 6, we present our results on the UCF101 dataset, training all methods using 4 A100 GPUs, with further scalability exploration conducted using 16 A100 GPUs. We mainly compare our method consistantly with Vision Mamba [123]. For the choice of the 3D Zigzag Mamba, please refer to Appendix 8.8. For Factorized 3D Zigzag Mamba in video processing, we deploy the sst scheme for factorizing spatial and temporal modeling. This scheme prioritizes spatial information complexity over temporal information, hypothesizing that redundancy exists in the temporal domain. Our results consistently demonstrate the superior performance of our method across various scenarios, underscoring the intricacy and effectiveness of our approach.
@@ 5 Conclusion
In this paper, we present the Zigzag Mamba Diffusion Model, developed within the Stochastic Interpolant framework. Our initial focus is on addressing the critical issue of spatial continuity. We then devise a Zigzag Mamba block with heterogeneous layerwise scan to better utilize the inductive bias in 2D images. Further, we factorize the 3D Mamba into 2D and 1D Zigzag Mamba to facilitate optimization. We empirically design various ablation studies to examine different factors. This approach allows for a more in-depth exploration of the Stochastic Interpolant theory. We hope our endeavor can inspire further exploration in the Mamba network design.
Acknowledgements
This project has been supported by the German Federal Ministry for Economic Affairs and Climate Action within the project “NXT GEN AI METHODS – Generative Methoden für Perzeption, Prädiktion und Planung”, the bidt project KLIMA-MEMES, Bayer AG, and the German Research Foundation (DFG) project 421703927. The authors gratefully acknowledge the Gauss Center for Supercomputing for providing compute through the NIC on JUWELS at JSC and the HPC resources supplied by the Erlangen National High Performance Computing Center (NHR@FAU funded by DFG).